OpenStudy (anonymous):

given NO + O3 --> NO2 + 02 what is the maximum volume of NO2 produced from 5.4mg NO and 9.6mg O3?

OpenStudy (anonymous):

@Kystal how do i know which element to use to find the max?

sam (.sam.):

You need to find which one is limiting reactant, either NO or O3

OpenStudy (anonymous):

first you find the n of moles of NO and O3~

OpenStudy (anonymous):

the from the equation we know that 1 mole of NO reacts with one mole of O3~ so the n of moles should be in the same amount~ but from the mass we know that one will be greater than another one~ the one with the less number of moles is referred as the limiting reagent~

OpenStudy (australopithecus):

Isn't this a limiting factor question

OpenStudy (anonymous):

@.Sam. okay limiting reagant: NO gives us 1.8x10-4mol of NO2 and O3 gives us 2x10^-4mol NO2, so its the one with the least amount NO2 produced?

OpenStudy (australopithecus):

I mean limiting reagent

OpenStudy (australopithecus):

then NO is the limiting reagent

OpenStudy (anonymous):

@Kystal so our limiting reagent is NO

OpenStudy (australopithecus):

if you did your calculations correctly then of course it is, limiting reagent means that it is the reactant that limits the amount of compound that can be made

OpenStudy (anonymous):

yeap~ NO is the limiting reagent~

sam (.sam.):

@bronzegoddess if your mole values are correct then yes

OpenStudy (anonymous):

okay, thank you.. for max values look for limiting reagent and for minimum as well?

OpenStudy (australopithecus):

The limiting reagent can be noted as the compound you need the most of to make the product

OpenStudy (australopithecus):

what ever makes the least amount of product will be the limiting reagent.

sam (.sam.):

\[1.8 \times 10^{-4} \text{mol of NO2}\times \frac{22.4 ~dm^{-3}~of NO_2}{1mol~of~NO_2}\]

OpenStudy (anonymous):

so now from the equation~ we know that n(NO)=n(NO2) =1.799x10^-4 mol n(NO2)=V/22.41 (using the formula n=V/22.41 at SPT) so V(NO2)=1.799x10^-4 (22.41) =4.03x10^-3 L (or 4.03 mL)